# Plug-and-Play Dramaturge: A Divide-and-Conquer Approach for Iterative Narrative Script Refinement via Collaborative LLM Agents

### v1

@@ Abstract
Figure 1: Our Dramaturge is inspired by the human scriptwriting process and performs Global Review, Scene-level Review, and Hierarchical Coordinated Revision to iteratively refine narrative scripts via a task and feature oriented divide-and-conquer strategy.
Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in generating diverse and engaging stories (Mirowski et al. 2023; Yang et al. 2023; Li et al. 2024; Zhao et al. 2024; Feng et al. 2025). However, existing LLM-based methods generate narratives from scratch using a single-pass workflow, even when multiple attempts are applied during generation. This overlooks a fundamental characteristic of creative practice: human scriptwriters craft scripts through multiple cycles of review and revision. As a result, current methods often struggle to produce high-quality scripts with coherent and nuanced storytelling(Tian et al. 2024a; Marco et al. 2024). This discrepancy highlights the importance of script refinement—a critical yet underexplored area in LLM-based creative writing, especially for long-form narrative generation (Tian et al. 2024b; Zhang et al. 2025).
How to effectively revise and improve long narrative scripts using LLMs remains a significant challenge because (Wu et al. 2025b, a; Chen et al. 2025):
@@ Introduction
To address these challenges, we propose Dramaturge, a divide-and-conquer approach for iterative narrative script refinement via collaborative LLM agents. This approach emulates humans’ workflow of initial reading, detailed reading, and in-depth revision, as illustrated in Figure 1. The strategy is applied across task and feature dimensions. First, the complex task is decomposed into three hierarchical stages: (1) a Global Review stage that performs a holistic analysis to grasp the storyline and identify structural issues; (2) a Scene-level Review stage that pinpoints local flaws and generates concrete, actionable suggestions for each scene; (3) a Hierarchical Coordinated Revision stage that revises the script coordinately across multiple locations. Second, through feature decomposition, each stage deploys a team of specialized agents, each focusing on a distinct dimension. The individual suggestions are consolidated to resolve conflicts and are then routed to the relevant agents in subsequent stages. The top-down task flow ensures that high-level strategies from the Global Review guide the detailed suggestions during the Scene-level Review, both of which jointly direct the local edits in the revision stage. Such coordination ensures that local modifications are aligned with the global narrative objectives. Finally, the entire workflow follows coarse-to-fine iterations by multiple rounds until the script quality stabilizes. This prevents superficial revisions by first adjusting the fundamental storyline before refining details, thereby achieving deep and comprehensive script enhancements.
@@
  • Comprehensive experiments show that Dramaturge significantly outperforms all strong baselines, improving upon the original scripts by 53.4% in script-level overall quality and 66.7% in scene-level details, and surpassing the strongest baseline by 8.3% and 19.9%, respectively.
Figure 2: The Architecture of Dramaturge. A task and feature oriented divide-and-conquer strategy is adopted, leveraging collaborative LLM agents to perform coarse-to-fine iterative refinement, while coordinating local edits and global structural adjustments and maintaining contextual consistency.
Related Work
LLM-Based Story Generation
@@
Narrative generation methods have gradually evolved from early generative neural networks (Fan, Lewis, and Dauphin 2018; Yao et al. 2019; Fan, Lewis, and Dauphin 2019; Yu et al. 2020; Ammanabrolu et al. 2021) to LLM-based approaches that leverage prompt engineering, hierarchical generation frameworks, and multi-agent role-playing. Dramatron (Mirowski et al. 2023) proposes a hierarchical generation framework that decomposes screenwriting into structured stages and generates high-quality long-form scripts. DOC (Yang et al. 2023) introduces detailed outline control to solve long-story coherence degradation problems. HoLLMwood (Chen et al. 2024b) adopts a role-playing mechanism in which specialized LLMs act as writers, editors, and actors, collaborating through iterative feedback cycles. IBSEN (Han et al. 2024) considers real-time user participation and proposes a director–actor collaboration framework to enable controllable script generation in human–AI interaction. FilmAgent (Xu et al. 2024) expands beyond traditional scriptwriting to encompass film production automation through multi-agent coordination. Ex3 (Lei et al. 2024) improves long-form novel generation by combining structure-aware instruction tuning with a tree-based expansion method, enhancing narrative coherence and character development. Neeko (Yu et al. 2024) adopts character-specific LoRA adapters and a gating mechanism to enable scalable multi-character role-playing. CRITICS (Bae and Kim 2024) enhances long-form story generation by incorporating a collaborative multi-agent critique process, addressing limitations in creativity and expressiveness of traditional plan-based methods. TaleForge (Nguyen et al. 2025) introduces an interactive multimodal system that enables personalized story creation through user-driven character and environment customization, addressing the growing demand for adaptive storytelling experiences. A knowledge-guided storytelling approach (Pan et al. 2025) leverages structured graphs to enhance narrative consistency and factual accuracy. R2 (Lin et al. 2025) proposes a Reader–Rewriter framework for adapting novels into screenplays, addressing hallucination and causality issues via refinement and plot graph construction. Domain-specific approaches (Cao et al. 2025) address cultural requirements by generating traditional opera scripts that preserve authenticity and narrative conventions. BookWorld (Ran et al. 2025) constructs multi-agent societies grounded in established fictional universes, addressing the challenge of simulating dynamic characters and consistent worldviews in book-based settings.
@@
Despite multiple attempts during generation, existing story generation methods still adhere to a single-pass workflow and lack iterative revisions employed by scriptwriters to progressively improve script quality. In this paper, we propose a pluggable framework that iteratively refines scripts in a coarse-to-fine manner, enabling systematic improvements while preserving contextual consistency.
LLM-driven Multi-agent Collaboration
@@
Recent work shows that collaborative LLM agents exhibit superior cognitive and reasoning capabilities over individual LLMs in complex tasks, and multi-agent collaboration has been widely applied to content creation (Zhang et al. 2024b; Durante et al. 2024; Guo et al. 2024; Xi et al. 2025; Qian et al. 2025). CAMEL (Li et al. 2023) proposes a communicative agent framework to study collaborative behaviors and multi-agent interactions. ChatDev (Qian et al. 2024) creates virtual companies where agents play professional roles to solve complex tasks through collaboration. AutoAgents (Chen et al. 2024a) develops an automatic agent generation framework to dynamically create and coordinate specialized agents based on task requirements. STORM (Shao et al. 2024) introduces a multi-agent writing system that synthesizes topic outlines through retrieval and multi-perspective question asking to generate long-form articles. Agents’ Room (Huot et al. 2025) consider story generation as a multi-step collaboration problem, proposing planning and writing agents to achieve enhanced story coherence. Chain of Agents (Zhang et al. 2024a) addresses long-context processing challenges by distributing tasks across multiple LLM agents in a sequential chain for improved efficiency and accuracy. Du et al. (Du et al. 2023) show that multi-agent debate mechanisms enhance factuality and reasoning capabilities in language models through collaborative deliberation processes. A multi-agent framework (Du et al. 2023) enhances factuality and reasoning in language models via collaborative deliberation. Debate-to-Write (Hu et al. 2025) introduces persona-driven multi-agent frameworks for diverse argument generation, enabling more comprehensive perspective coverage. AgentCoord (Pan et al. 2024) provides visual exploration tools for understanding and optimizing coordination strategies in LLM-based multi-agent systems.
@@
Existing work primarily adopts role-playing and collaboration to simplify complex tasks; however, in our task, it remains challenging to coordinate local modifications with the global narrative. We leverage the logical reasoning power of multi-agent LLMs to implement divide-and-conquer strategies along both task and feature dimensions, enabling coordinated revisions across multiple granularities and locations.
[…]
Output: Refined script SS
[…]
3:  // Global Review
[…]
7:  // Hierarchical Coordinated Revision - structural refinement
[…]
10:  end if
11:  // Scene-level Review
[…]
15:  end for
[…]
17:  // Hierarchical Coordinated Revision - detail refinement
[…]
20:  end for
[…]
22:  // Evaluating the quality improvement
[…]
26:  end if
[…]
28:   break
29:  end if
30:  ii+1$\leftarrow$ i+1
31:end while
[…]
Methodology
Overall Architecture
@@
The system architecture is illustrated in Figure 2. We adopt a divide-and-conquer strategy along task and feature dimensions. For task decomposition, a three-stage hierarchical structure is employed: (1) Global Review performs holistic analysis based on the full script and summary to formulate high-level improvement strategies; (2) Scene-level Review traverses each scene and produces detailed modification suggestions following global strategies; (3) Hierarchical Coordinated Revision coordinates modifications across multiple granularities and locations. For feature decomposition, each stage deploys specialized agents targeting distinct review dimensions. Their suggestions are integrated to resolve conflicts and redundancies, then routed to subsequent stages. The refinement workflow proceeds iteratively in a coarse-to-fine manner across multiple rounds until no substantive improvement can be made. The algorithmic procedure is presented in Algorithm 1.
Global Review
@@
The Global Review Integrator consolidates suggestions from the four evaluators by resolving potential conflicts and synthesizing them into coherent improvement strategies aligned with the overall narrative structure. The integrated guidance is then distributed to the Storyline Editor for main plot modifications and the Scene-level Review for detailed scene-specific inspection.
@@
Scene-level Review
The Scene-level Review (SLR) traverses each scene to generate concrete, scene-level modification suggestions under the guidance of the GR. After that, the results are also synthesized into component-specific guidance for subsequent revision, ensuring coherence and precision in the process.
@@
Following the inspection process, the Scene-level Review Integrator organizes the refined suggestions into a unified whole by identifying and eliminating overlapping content, creating actionable improvement directives tailored to each scene’s specific context. Subsequently, the Suggestion Router classifies each modification by its intended revision domain and routes it to either the subsequent Dialogue Editor or the Scene Editor accordingly. This systematic routing maintains the specialized focus of each editor while avoiding conflicting modification. The prompts for the specialized agents in the SLR stage are provided in Appendix B.
Table 1: Quantitative Evaluation Results. Script-level scores evaluate the complete scripts, while scene-level scores are aggregated from scene-by-scene comparisons, emphasizing the quality of details. Despite using a relatively weaker backbone, our Dramaturge significantly outperforms the original scripts and six strong baselines in both overall quality and details. Bold indicates significance with p<0.05p<0.05.
Original Scripts 12.94 15.12 13.02 16.10 57.18 56.50
GPT-4o 14.00 $\uparrow$8.215.96 $\uparrow$5.613.72 $\uparrow$5.416.98 $\uparrow$5.560.66 $\uparrow$6.154.12 $\downarrow$4.2%
Qwen-max 13.78 $\uparrow$6.515.90 $\uparrow$5.213.76 $\uparrow$5.716.82 $\uparrow$4.560.26 $\uparrow$5.454.28 $\downarrow$3.9%
Deepseek-r1 16.56 $\uparrow$28.017.16 $\uparrow$13.514.98 $\uparrow$15.120.04 $\uparrow$24.568.74 $\uparrow$20.266.24 $\uparrow$17.2%
GPT-4.1-mini 16.94 $\uparrow$30.918.42 $\uparrow$21.816.28 $\uparrow$25.019.64 $\uparrow$22.071.28 $\uparrow$24.763.92 $\uparrow$13.1%
GPT-4.1 17.38 $\uparrow$34.318.76 $\uparrow$24.117.10 $\uparrow$31.320.10 $\uparrow$24.873.34 $\uparrow$28.365.00 $\uparrow$15.0%
Gemini-2.5-pro 19.28 $\uparrow$49.021.00 $\uparrow$38.918.58 $\uparrow$42.722.12 $\uparrow$37.480.98 $\uparrow$41.678.56 $\uparrow$39.1%
@@
Hierarchical Coordinated Revision
@@
Iterative Improvement with Quality Control
@@
The system transitions from structural refinement to detail refinement when no substantive improvement can be made, and the iterative process terminates if detail refinement fails to yield improvement within N attempts. A substantive improvement is defined as an increase of at least 1.0 point in the Script-Level Overall Evaluation, avoiding unnecessary and frequent revisions caused by minor score fluctuations.
Experiments
Experimental Settings
Implementation Details
Our Dramaturge uses GPT-4.1-mini as the backbone model with temperature = 0.7 and top-p = 0.95 for script refinement to achieve better creative outputs. We vary the number of script revision retries from 1 to 10. Performance plateaus when N>3, so we set the maximum retry count to N=3 for efficiency and cost considerations. For evaluation, we set temperature = 0 to obtain more stable and consistent evaluation results.
Dataset
To facilitate a comprehensive and balanced evaluation, we constructed a dataset of 50 narrative scripts across five genres, evenly sourced from five representative works, with a mix of human-written and AI-generated scripts. The dataset is constructed from the following work:
  • Writingprompts (Fan, Lewis, and Dauphin 2018): A collection of creative stories from the Reddit community.
  • Dramatron (Mirowski et al. 2023): High-quality generated scripts using a hierarchical generation strategy.
  • DOC (Yang et al. 2023): Long stories generated using a two-stage outline control strategy.
  • MoPS (Ma, Qiao, and Liu 2024): Generated scripts leveraging Well-crafted modular premise synthesis.
  • Agents’ Room (Huot et al. 2025): High-quality stories collected through collaborative writing.

### v2

@@ Abstract.
Narrative generation, storytelling, parallel art, large language models, LLM agents
copyright: none
Figure 1. Our Dramaturge is inspired by the human scriptwriting process and performs Global Review, Scene-level Review, and Hierarchical Coordinated Revision to iteratively refine narrative scripts via a task and feature oriented divide-and-conquer strategy.
1. Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in generating diverse and engaging stories (Mirowski et al., 2023; Yang et al., 2023; Li et al., 2024; Zhao et al., 2024; Feng et al., 2025).
However, existing LLM-based methods generate narratives from scratch using a single-pass workflow, even when multiple attempts are applied during generation. This overlooks a fundamental characteristic of creative practice: human scriptwriters craft scripts through multiple cycles of review and revision (Nelmes, 2008; Batty and Taylor, 2021) to progressively improve script quality. As a result, current methods often struggle to produce high-quality scripts with coherent and nuanced storytelling (Tian et al., 2024; Marco et al., 2024). This discrepancy highlights the importance of script refinement—a critical yet underexplored area in LLM-based creative writing, especially for long-form narrative generation (Gervás et al., 2004; Mulholland et al., 2015; Porteous et al., 2010; Guo et al., 2021; Lisena et al., 2023; Zhang et al., 2025).
How to effectively revise and improve long narrative scripts using LLMs remains a significant challenge because (Wu et al., 2025b, a; Chen et al., 2025):
@@ 1. Introduction
To address these challenges, we propose Dramaturge, a divide-and-conquer approach for iterative narrative script refinement via collaborative LLM agents. This approach emulates scriptwriters’ workflow of initial reading, detailed reading, and in-depth revision, as illustrated in Fig. 1. The strategy is applied across task and feature dimensions. First, the complex task is decomposed into three hierarchical stages: (1) a Global Review stage that performs a holistic analysis to grasp the storyline and identify structural issues; (2) a Scene-level Review stage that pinpoints local flaws and generates concrete, actionable suggestions for each scene; (3) a Hierarchical Coordinated Revision stage that revises the script coordinately across multiple locations. Second, through feature decomposition, each stage deploys a team of specialized agents, each focusing on a distinct dimension. The individual suggestions are consolidated to resolve conflicts and are then routed to the relevant agents in subsequent stages. The top-down task flow ensures that high-level strategies from the Global Review guide the detailed suggestions during the Scene-level Review, both of which jointly direct the local edits in the revision stage. Such coordination ensures that local modifications are aligned with the global narrative objectives. Finally, the entire workflow follows coarse-to-fine iterations by multiple rounds until the script quality stabilizes. This prevents superficial revisions by first adjusting the fundamental storyline before refining details, thereby achieving deep and comprehensive script enhancements.
@@
  • Comprehensive experiments show that Dramaturge significantly outperforms all strong baselines, improving upon the original scripts by 53.2% in script-level overall quality and 65.1% in scene-level details, and surpassing the strongest baseline by 8.2% and 18.4%, respectively.
2. Related Work
2.1. LLM-Based Narrative Generation
@@
Narrative generation methods have gradually evolved from early generative neural networks (Fan et al., 2018; Yao et al., 2019; Fan et al., 2019; Yu et al., 2020; Ammanabrolu et al., 2021) to LLM-based approaches that leverage prompt engineering, hierarchical generation frameworks, and multi-agent role-playing. Dramatron (Mirowski et al., 2023) proposes a hierarchical generation framework that decomposes screenwriting into structured stages and generates high-quality long-form scripts. DOC (Yang et al., 2023) introduces detailed outline control to solve long-story coherence degradation problems. HoLLMwood (Chen et al., 2024b) adopts a role-playing mechanism in which specialized LLMs act as writers, editors, and actors, collaborating through iterative feedback cycles. IBSEN (Han et al., 2024) considers real-time user participation and proposes a director–actor collaboration framework to enable controllable script generation in human–AI interaction. FilmAgent (Xu et al., 2024) expands beyond traditional scriptwriting to encompass film production automation through multi-agent coordination. Ex3 (Lei et al., 2024) improves long-form novel generation by combining structure-aware instruction tuning with a tree-based expansion method, enhancing narrative coherence and character development. Neeko (Yu et al., 2024) adopts character-specific LoRA adapters and a gating mechanism to enable scalable multi-character role-playing. CRITICS (Bae and Kim, 2024) enhances long-form story generation by incorporating a collaborative multi-agent critique process, addressing limitations in creativity and expressiveness of traditional plan-based methods. TaleForge (Nguyen et al., 2025) introduces an interactive multimodal system that enables personalized story creation through user-driven character and environment customization, addressing the growing demand for adaptive storytelling experiences. A knowledge-guided storytelling approach (Pan et al., 2025) leverages structured graphs to enhance narrative consistency and factual accuracy. R2 (Lin et al., 2025) proposes a Reader–Rewriter framework for adapting novels into screenplays, addressing hallucination and causality issues via refinement and plot graph construction. Domain-specific approaches (Cao et al., 2025) address cultural requirements by generating traditional opera scripts that preserve authenticity and narrative conventions. BookWorld (Ran et al., 2025) constructs multi-agent societies grounded in established fictional universes, addressing the challenge of simulating dynamic characters and consistent worldviews in book-based settings.
@@
Despite multiple attempts during generation, existing narrative generation methods still adhere to a single-pass workflow and lack iterative revisions employed by scriptwriters to progressively improve script quality. This limitation constrains the quality of the generated scripts. In this paper, we propose a pluggable framework that iteratively refines scripts in a coarse-to-fine manner, enabling systematic improvements while preserving contextual consistency.
2.2. LLM-driven Multi-agent Systems
@@
Recent work shows that collaborative LLM agents exhibit superior cognitive and reasoning capabilities over individual LLMs in complex tasks, and multi-agent collaboration has been widely applied to content creation (Zhang et al., 2024b; Durante et al., 2024; Guo et al., 2024; Xi et al., 2025; Qian et al., 2025). CAMEL (Li et al., 2023) proposes a communicative agent framework to study collaborative behaviors and multi-agent interactions. ChatDev (Qian et al., 2024) creates virtual companies where agents play professional roles to solve complex tasks through collaboration. AutoAgents (Chen et al., 2024a) develops an automatic agent generation framework to dynamically create and coordinate specialized agents based on task requirements. STORM (Shao et al., 2024) introduces a multi-agent writing system that synthesizes topic outlines through retrieval and multi-perspective question asking to generate long-form articles. Agents’ Room (Huot et al., 2025) consider story generation as a multi-step collaboration problem, proposing planning and writing agents to achieve enhanced story coherence. Chain of Agents (Zhang et al., 2024a) addresses long-context processing challenges by distributing tasks across multiple LLM agents in a sequential chain for improved efficiency and accuracy. A multi-agent framework (Du et al., 2023) enhances factuality and reasoning in language models via collaborative deliberation. Debate-to-Write (Hu et al., 2025) introduces persona-driven multi-agent frameworks for diverse argument generation, enabling more comprehensive perspective coverage. AgentCoord (Pan et al., 2024) provides visual exploration tools for understanding and optimizing coordination strategies in LLM-based multi-agent systems.
@@
LLM-driven multi-agent systems are widely used to decompose and tackle complex tasks. In our task, coordinating local modifications with global narrative coherence remains challenging. We leverage the logical reasoning capabilities of LLM agentic systems and propose divide-and-conquer strategies along both task and feature dimensions to revise long-context scripts, enabling coordinated revisions across multiple granularities and locations under global guidance.
2.3. Parallel Art
Beyond narrative generation, broader frameworks for AI-assisted creation have been proposed. Inspired by parallel systems and ACP methods(Wang, 2004b, a, 2007; Wang et al., 2016), Parallel Art (Guo et al., 2019, 2023b, 2023a) establishes a framework for human-machine collaborative artistic creation through computational experiments and iterative optimization between virtual and actual systems. Building on this foundation, Parallel Theaters (Ni et al., 2023, 2025) further introduces a drama-creation approach that coordinates human, digital, and robotic artists. Through narrative reasoning and progression, it transforms decision-making processes into immersive dramatic interactions, thereby enabling what is known as a decision theater.
Following the notion of computational experiments, we build a closed-loop pipeline for generation, evaluation, and optimization, enabling iterative revision and progressive improvement of long-form narratives.
3. Methodology
3.1. Overall Architecture
@@
The system architecture is illustrated in Fig. 2. We adopt a divide-and-conquer strategy along task and feature dimensions. For task decomposition, a three-stage hierarchical structure is employed: (1) Global Review performs holistic analysis based on the full script and summary to formulate high-level improvement strategies; (2) Scene-level Review traverses each scene and produces detailed modification suggestions following global strategies; (3) Hierarchical Coordinated Revision coordinates modifications across multiple granularities and locations. For feature decomposition, each stage deploys specialized agents targeting distinct review dimensions. Their suggestions are integrated to resolve conflicts and redundancies, then routed to subsequent stages. The refinement workflow proceeds iteratively in a coarse-to-fine manner across multiple rounds until no substantive improvement can be made. The algorithmic procedure is presented in Algorithm 1.
3.2. Global Review
@@
Figure 2. The Architecture of Dramaturge. A task and feature oriented divide-and-conquer strategy is adopted, leveraging collaborative LLM agents to perform coarse-to-fine iterative refinement, while coordinating local edits and global structural adjustments and maintaining contextual consistency.
The Global Review Integrator consolidates suggestions from four evaluators prioritizing each suggestion according to both scope and magnitude of its impact. We first group suggestions by scope: (1) structural directives that reshape narrative arcs, character development, or causal relations spanning multiple scenes; (2) cross-scene adjustments that affect narrative continuity or thematic resonance; (3) localized detail optimizations. Items in a higher scope group always outrank those in lower groups. Within each group, suggestions tightly coupled with the main plotline take precedence over secondary subplots, and cross-scene effects surpass single-scene ones. When two items still tie, we follow the order: Narrative, Character, Theme, and Engagement. Higher-priority directives are retained, with compatible lower-priority ones merged when possible and conflicting ones discarded. The resulting non-overlapping plan routes plot-critical guidance to the Storyline Editor, while remaining directives flow to the Scene-level Review for fine-grained inspection.
@@
3.3. Scene-level Review
The Scene-level Review (SLR) traverses each scene to generate concrete, scene-level modification suggestions under the guidance of the GR, prioritizing global consistency. When a scene-level recommendation conflicts with GR directives, the local suggestion is subordinated to the global narrative strategy to eliminate potential inconsistencies. The resulting suggestions are then synthesized into component-specific guidance for subsequent revision, ensuring coherence and precision in the process.
@@ 3.3. Scene-level Review
Following the inspection, the Scene-level Review Integrator scores each recommendation with a three-stage priority policy to resolve conflicts: (1) compliance with directives from the Global Review is mandatory, with conflicting local suggestions being immediately rejected and alignments with GR guidance receiving top priority; (2) within the same inspector tier, we favor interventions whose impact propagates to subsequent scenes (e.g., foreshadowing, metaphors) or multiple dialogue exchanges ahead of purely stylistic tweaks; (3) among the remaining candidates, inspector type dictates precedence: Plot and Character Inspectors rank above Dialogue and Scene Description Inspectors because they most directly influence causality and behavioral authenticity. Higher-priority items are preserved, compatible lower-priority ones are merged, and incompatible ones are dropped. The resulting directives are mapped to specific dialogue or descriptive spans to provide actionable, context-aware guidance for the subsequent editors.
Subsequently, the Suggestion Router classifies each modification by its intended revision domain and routes it to either the subsequent Dialogue Editor or the Scene Editor accordingly. This systematic routing maintains the specialized focus of each editor while avoiding conflicting modifications. The prompts for the specialized agents in the SLR stage are provided in Appendix B.
@@
2:Output: Refined script SS
[…]
5:  // Global Review
[…]
9:  // Hierarchical Coordinated Revision - structural refinement
[…]
12:  end if
13:  // Scene-level Review
[…]
17:  end for
[…]
19:  // Hierarchical Coordinated Revision - detail refinement
[…]
23:  end for
[…]
26:  // Evaluating the quality improvement
[…]
30:  end if
[…]
32:   break
33:  end if
[…]
35:  ii+1$\leftarrow$ i+1
36:end while
[…]
3.4. Hierarchical Coordinated Revision
@@
3.5. Iterative Improvement with Quality Control
@@
[…]
4. Experiments
4.1. Experimental Settings
4.1.1. Implementation Details
[…]
For evaluation, we employ Gemini-2.5-pro with temperature = 0 to obtain stable and rigorous assessment results. This setting is grounded in our comparative analysis of mainstream flagship models from the Gemini, GPT, Qwen, and DeepSeek series. Pairwise Kolmogorov–Smirnov (KS) tests on the evaluation outputs revealed no significant differences in score distributions across models (p>0.1p>0.1), indicating broad consensus in their evaluation standards. However, among these statistically consistent models, Gemini-2.5-pro exhibited the most stringent absolute scoring. Therefore, we adopt it as a strict evaluator to enforce a high-quality standard. Furthermore, using Gemini as the evaluation model mitigates potential self-preference bias that may arise when the same model family (GPT) is used for both generation and evaluation.
Table 1. Quantitative Evaluation Results. Script-level scores evaluate the complete scripts, while scene-level scores are aggregated from scene-by-scene comparisons, emphasizing the quality of details. Despite using a relatively weaker backbone, our Dramaturge significantly outperforms the original scripts and six strong baselines in both overall quality and details. Bold indicates significance with p<0.05p<0.05.
Original Scripts 12.50 14.60 12.58 15.56 55.24 54.50
GPT-4o 13.52 $\uparrow$8.215.42 $\uparrow$5.613.24 $\uparrow$5.216.40 $\uparrow$5.458.58 $\uparrow$6.052.20 $\downarrow$4.2%
Qwen-max 13.30 $\uparrow$6.415.36 $\uparrow$5.213.28 $\uparrow$5.616.24 $\uparrow$4.458.18 $\uparrow$5.352.36 $\downarrow$3.9%
Deepseek-r1 16.00 $\uparrow$28.016.56 $\uparrow$13.414.46 $\uparrow$14.919.36 $\uparrow$24.466.38 $\uparrow$20.263.96 $\uparrow$17.4%
GPT-4.1-mini 16.36 $\uparrow$30.917.78 $\uparrow$21.815.72 $\uparrow$25.018.96 $\uparrow$21.968.82 $\uparrow$24.661.70 $\uparrow$13.2%
GPT-4.1 16.78 $\uparrow$34.218.10 $\uparrow$24.016.50 $\uparrow$31.219.40 $\uparrow$24.770.78 $\uparrow$28.162.72 $\uparrow$15.1%
Gemini-2.5-pro 18.60 $\uparrow$48.820.28 $\uparrow$38.917.94 $\uparrow$42.621.36 $\uparrow$37.378.18 $\uparrow$41.576.00 $\uparrow$39.4%
[…]
4.1.2. Dataset
Due to the substantial length of individual scripts and the significant computational overhead involved in their processing and evaluation, previous studies typically used datasets with only a few dozen samples (e.g., DOC (Yang et al., 2023), 20 scripts; HoLLMwood (Chen et al., 2024b), 60 scripts). Thus, to achieve broader coverage while remaining within cost and efficiency constraints, we construct a dataset of 100 narrative scripts drawn evenly from five sources:
  • Writingprompts (Fan et al., 2018): A collection of creative stories from the Reddit community.
  • Dramatron (Mirowski et al., 2023): High-quality generated scripts using a hierarchical generation strategy.
  • DOC (Yang et al., 2023): Long stories generated using a two-stage outline control strategy.
  • MoPS (Ma et al., 2024): Generated scripts leveraging Well-crafted modular premise synthesis.
  • Agents’ Room (Huot et al., 2025): High-quality stories collected through collaborative writing.